(4%) Problem! 11: A 200 turn circular loop of radius 65 cmis lying flat on the ground with a curent of 95 A circulating counterclockwise (when viewed from above) in a location where the Earths field is pointed north, but at an angle 45.0° below the horizontal and with a strength of 5.85 x 10 T What is the magnitude of the torque on the loop, in N m? Grade Summ Potential sin0 cotano asin0 Submissions Attempts rema (1% per attem detailed view coso tano acoso atan acotan sinh) cosh tanh cotanh) O Degrees O Radians I give up! Hints:--deduction per hint. Hints remaining:- Feedback: 0% deduction per feedback.
Answer #1

here,

radius , r = 65 cm = 0.65 m

magnetic feild , B = 5.85 * 10^-5 T

theta = 45 degree

current , I = 95 A

the magnitude of torque on the loop , T = N * I * A * B * sin(theta)

T = 1 * 95 * pi * 0.65^2 * 5.85 * 10^-5 * sin(45) N.m

T = 5.21 * 10^-3 N.m

the magnitude of torque on the loop is 5.21 * 10^-3 N.m